I'm having trouble solving the Logarithmic equation Log(Base 2)(x-5)=4 . So far I've tried Moving the 4 over to the 2 to make 2^4=x-5, but after that I'm stuck. Could someone show me a step by step way to achieve the answer? Thanks!
Found 2 solutions by lwsshak3, nerdybill:
Answer by lwsshak3(11628)

I'm having trouble solving the Logarithmic equation Log(Base 2)(x-5)=4 . So far I've tried Moving the 4 over to the 2 to make 2^4=x-5, but after that I'm stuck. Could someone show me a step by step way to achieve the answer?
**
log 2(x-5)=4
By definition, this says the logarithm of (x-5), base 2 is equal to 4, that is, the base(2) raised to the logarithm of the number(4) is equal to the number(x-5)
Thus, 2^4=x-5
16=x-5
x=21
Check:
log2(x-5)=log2(21-5)=log2(16)=4

Answer by nerdybill(7384)